can 10, 20 and 30 can be the lenghts of the sides of the triangle? right triangle? Explain
Answer by dabanfield(803)

If the triangle is a right triangle then the hypotenuse (the longest side) is 30. From the Pythagorean Theorem we would then have 10^2 + 20^2 = 30^2 or
100 + 400 = 900
Since 500 does not equal 900 we must not have a right triangle.